OpenStudy (anonymous):

HELPP ! a country's population in 1993 was 171 million. in 1999 it was 176 million. estimate the population in 2012 using the exponential growth formula

OpenStudy (anonymous):

??

OpenStudy (fifciol):

\[x_t=x_0(1+r)^t\] \[176=171(1+r)^{1999-1993} \rightarrow r=(\frac{ 176 }{171 })^{\frac{ 1 }{ 6 }}\] \[x=176*(\frac{ 176 }{171 })^{\frac{ 13 }{ 6 }}= 187,3 million\]